Use the tabulated half-cell potentials to calculate the equilibrium constant (K) for the following balanced redox reaction at 25
Answer #1

3I_2+6e^-\rightarrow 6I^-   E^o(reduction/cathode)= 0.54V

2Fe\rightarrow 2Fe^3^++6e^-E^o(oxidation/anode)=0.04V

E^o=E^o(reduction)+E^o(oxidation)

E^o=0.58V

E^o=\frac{RT}{nF}lnK

E^o=0.58V

R=8.314\frac{J}{molK}

T=298K

n=6e^-

F=96485\frac{Coulomb}{mole}

lnK=135.52

K=7.19*10^{58}

K\approx 2.4*10^5^8
Answer #2

1.1
Answer #3
1.7